# Conditional PED-ANOVA: Hyperparameter Importance in Hierarchical & Dynamic Search Spaces

### v1

@@ D.2. Implementation Details for Baselines
Appendix E Supplementary Analysis for the Ablation Study Using Standard Local HPI
E.1. Leakage of Gating Effects in Standard Local HPI
@@
Let $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{\bm{x}\in\mathcal{X}_{\gamma^{\prime}}\}$ be the top-set indicator. Consider a gating hyperparameter $x^{(c)}$ and a conditional (child) hyperparameter $x^{(d)}$ whose regime is determined by $x^{(c)}$. Formally, assume that for a conditional hyperparameter $x^{(d)}$ there exists a measurable mapping $\phi^{(d)}\colon\mathcal{X}^{(c)}\to\{1,\ldots,K^{(d)}\}$ such that:
@@
Here, $x^{(c)}$ corresponds to the gating variable $c$ in the synthetic objectives (Equations 6 and 9), while $x^{(d)}$ corresponds to the conditional variables $x$ and $y$. These assumptions are satisfied by the synthetic objectives used in our experiments.
@@
Theorem E.1 (Leakage of gating effects under the standard local HPI).
Under the above assumptions (Equations 17 and 18), the standard local marginal variance of the conditional hyperparameter $x^{(d)}$ contains the local marginal variance of the gating hyperparameter $x^{(c)}$:
@@
The proof is provided in Section E.2. Consequently, when HPI is computed from $v_{\gamma}^{(d)}$, the effect of the gating variable $x^{(c)}$ is additively inherited by conditional hyperparameters through the inter-regime term.
This theorem is consistent with our empirical observations. In Figure 3, when using the standard local HPI, the importance of the gating variable $c$ is always no larger than those of $x$ and $y$. Moreover, in the regions of $\gamma^{\prime}$ where $x$ or $y$ is inactive, their HPI becomes almost identical to that of $c$. This is exactly what Theorem E.1 predicts: once $v_{\gamma,\mathrm{within}}^{(d)}$ vanishes for an inactive variable, the standard local variance $v_{\gamma}^{(d)}$ reduces to the additive gating term $v_{\gamma}^{(c)}$.
E.2. Proof of Theorem E.1
@@
$=\mathbb{E}\!\left[b_{\gamma^{\prime}}\mid I^{(d)}\right].$
@@
$f_{\gamma}^{(c)}(x^{(c)})$ $\coloneqq\frac{1}{Z_{\gamma}^{(c)}(x^{(c)})}\sum_{n=1}^{N}b_{\gamma^{\prime}}(x^{(c)},x^{(\bar{c}\,)}_{n})\,b_{\gamma}(x^{(c)},x^{(\bar{c}\,)}_{n})$,
@@
Comparing Equations 21, 22 and 23, we conclude that the inter-regime variance of the conditional hyperparameter $x^{(d)}$ equals the marginal variance of the gating hyperparameter $x^{(c)}$:
$v_{\gamma,\mathrm{inter}}^{(d)}=v_{\gamma}^{(c)}.$
@@
E.3. Decomposition of PED over Regimes
@@
Lemma E.2 (Regime-wise decomposition of Pearson divergence).
@@
The proof is provided in Section E.4. The absolute continuity $\mu_{\gamma^{\prime}}^{(d)}\ll\mu_{\gamma}^{(d)}$ holds for the extended marginals in the same way as in Equation 1. The regime-wise $\mu_{\gamma^{\prime},\,i}^{(d)}\ll\mu_{\gamma,\,i}^{(d)}$ also holds for all regimes with $\beta_{i}^{(d)}>0$, since absolute continuity is preserved under restriction to measurable subsets. If a regime $i$ corresponds to an inactive configuration (i.e., $\mathcal{Z}_{i}^{(d)}=\{\bot\}$), then $\mu_{\gamma^{\prime},\,i}^{(d)}$ and $\mu_{\gamma,\,i}^{(d)}$ are both degenerate at $\bot$, so $D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,\mu_{\gamma,\,i}^{(d)}\right)=0$ and only the inter-regime term contributes.
@@
$\frac{d\mu_{\gamma^{\prime}}}{d\mu_{\gamma}}(x)=\frac{d(\alpha_{i}\,\mu_{\gamma^{\prime},\,i})}{d(\beta_{i}\,\mu_{\gamma,\,i})}(x)=\frac{\alpha_{i}}{\beta_{i}}\,\frac{d\mu_{\gamma^{\prime},\,i}}{d\mu_{\gamma,\,i}}(x),$
@@
By substituting Equations 26 and 27 into Equation 25, we have:
$D_{\mathrm{PE}}(\mu_{\gamma^{\prime}}\|\mu_{\gamma})$ $=\sum_{i=1}^{\;K^{(d)}}\left(\frac{\alpha_{i}^{2}}{\beta_{i}}\left(D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+1\right)-2\alpha_{i}+\beta_{i}\right)$
@@
This completes the proof.

### v2

@@ E.2.2. Results and Discussion
Appendix F Supplementary Analysis for the Ablation Study Using Standard Local HPI
In this section, we provide additional theoretical details for the ablation study on the effect of within-regime variance for HPI presented in Section 5.5.1. In Section F.1, we give a theoretical analysis showing that gating effects can leak when within-regime local marginal variance is not used. In addition, Section 5.5.1 compared condPED-ANOVA against a PED-ANOVA-style construction of the standard local HPI using $v_{\gamma}^{(d)}$ as an ablation baseline (Equation 17); we provide its derivation in Section F.3.
F.1. Leakage of Gating Effects in Standard Local HPI
@@
Let $b_{\gamma^{\prime}}\coloneqq\mathbf{1}\{\bm{x}\in\mathcal{X}_{\gamma^{\prime}}\}$ be the top-set indicator. Consider a conditioning hyperparameter $x^{(c)}$ and a conditioned hyperparameter $x^{(d)}$ whose regime is determined by $x^{(c)}$. Formally, assume that for a conditional hyperparameter $x^{(d)}$ there exists a measurable mapping $\phi^{(d)}\colon\mathcal{X}^{(c)}\to\{1,\ldots,K^{(d)}\}$ such that:
@@
Here, $x^{(c)}$ corresponds to the gating variable $c$ in the synthetic objectives (Equations 12 and 15), while $x^{(d)}$ corresponds to the conditioned variables $x$ and $y$. These assumptions are satisfied by the synthetic objectives used in our experiments.
@@
Theorem F.1 (Leakage of gating effects under the standard local HPI).
Under the above assumptions (Equations 36 and 37), the standard local marginal variance of the conditional hyperparameter $x^{(d)}$ contains the local marginal variance of the gating hyperparameter $x^{(c)}$:
@@
The proof is provided in Section F.2. Consequently, when HPI is computed from $v_{\gamma}^{(d)}$, the effect of the gating variable $x^{(c)}$ is additively inherited by conditional hyperparameters through the inter-regime term.
This theorem is consistent with our empirical observations. In Figure 3, when using the standard local HPI, the importance of the gating variable $c$ is always no larger than those of $x$ and $y$. Moreover, in the regions of $\gamma^{\prime}$ where $x$ or $y$ is inactive, their HPI becomes almost identical to that of $c$. This is exactly what Theorem F.1 predicts: once $v_{\gamma,\mathrm{within}}^{(d)}$ vanishes for an inactive variable, the standard local variance $v_{\gamma}^{(d)}$ reduces to the additive gating term $v_{\gamma}^{(c)}$.
F.2. Proof of Theorem F.1
@@
(40) $=\mathbb{E}\!\left[b_{\gamma^{\prime}}\mid I^{(d)}\right].$
@@
(42) $f_{\gamma}^{(c)}(x^{(c)})$ $\coloneqq\frac{1}{Z_{\gamma}^{(c)}(x^{(c)})}\sum_{n=1}^{N}b_{\gamma^{\prime}}(x^{(c)},x^{(\bar{c}\,)}_{n})\,b_{\gamma}(x^{(c)},x^{(\bar{c}\,)}_{n})$,
@@
Comparing Equations 41, 43 and 44, we conclude that the inter-regime variance of the conditional hyperparameter $x^{(d)}$ equals the marginal variance of the gating hyperparameter $x^{(c)}$:
(45) $v_{\gamma,\mathrm{inter}}^{(d)}=v_{\gamma}^{(c)}.$
@@
F.3. Decomposition of PED over Regimes
@@
Lemma F.2 (Regime-wise decomposition of Pearson divergence).
@@
The proof is provided in Section F.4. The absolute continuity $\mu_{\gamma^{\prime}}^{(d)}\ll\mu_{\gamma}^{(d)}$ holds for the extended marginals in the same way as in Equation 5. The regime-wise $\mu_{\gamma^{\prime},\,i}^{(d)}\ll\mu_{\gamma,\,i}^{(d)}$ also holds for all regimes with $\beta_{i}^{(d)}>0$, since absolute continuity is preserved under restriction to measurable subsets. If a regime $i$ corresponds to an inactive configuration (i.e., $\mathcal{Z}_{i}^{(d)}=\{\bot\}$), then $\mu_{\gamma^{\prime},\,i}^{(d)}$ and $\mu_{\gamma,\,i}^{(d)}$ are both degenerate at $\bot$, so $D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}^{(d)}\,\middle\|\,\mu_{\gamma,\,i}^{(d)}\right)=0$ and only the inter-regime term contributes.
@@
(47) $\frac{d\mu_{\gamma^{\prime}}}{d\mu_{\gamma}}(x)=\frac{d(\alpha_{i}\,\mu_{\gamma^{\prime},\,i})}{d(\beta_{i}\,\mu_{\gamma,\,i})}(x)=\frac{\alpha_{i}}{\beta_{i}}\,\frac{d\mu_{\gamma^{\prime},\,i}}{d\mu_{\gamma,\,i}}(x),$
@@
By substituting Equations 49 and 50 into Equation 48, we have:
(51) $D_{\mathrm{PE}}(\mu_{\gamma^{\prime}}\|\mu_{\gamma})$ $=\sum_{i=1}^{\;K^{(d)}}\left(\frac{\alpha_{i}^{2}}{\beta_{i}}\left(D_{\mathrm{PE}}\left(\mu_{\gamma^{\prime},\,i}\,\middle\|\,\mu_{\gamma,\,i}\right)+1\right)-2\alpha_{i}+\beta_{i}\right)$
@@
This completes the proof. ∎
Appendix G Additional Results and Discussion on Experiments on Real-World Benchmarks (Section 6)
In Section 6, we presented results on the YAHPO Gym rbv2_super scenario (instance ID 1053). Here, we report additional results on other instances of the rbv2_super benchmark suite provided by YAHPO Gym. The results are shown in Figure 14. The corresponding statistics of objective values (acc) for each learner_id are provided in Table 2.
These results show that, across instances, learner_id, train_size, and hyperparameters associated with the top-performing learners occupy the top ranks in HPI. Moreover, the HPIs produced by condPED-ANOVA broadly match the learner-wise performance ordering in Table 2, supporting their validity. In instance ID 1468, although aknn attains the highest mean performance, svm and ranger achieve higher maximum performance. This is reflected in higher HPIs in svm and ranger, which is reasonable given that our goal is to achieve the highest possible score.
In instance IDs 1063 and 1468, the performance gaps among the top learners (in terms of mean and maximum accuracy) are smaller than in instance 1457. Consistently, for these instances, the HPI of learner_id is slightly lower than that of hyperparameters associated with the top learners, which is reasonable.
Moreover, across all instances, condPED-ANOVA consistently assigns high HPI to empirically known important hyperparameters within each learner, such as sample.fraction and min.node.size for ranger, the cost and kernel for svm, and k for aknn. This is reasonable given their well-known influence on model performance.
These results demonstrate that condPED-ANOVA consistently yields meaningful and interpretable HPIs across multiple real-world instances of the rbv2_super suite, emphasizing its robustness and applicability in practical scenarios.
Table 2. Summary statistics of the objective value (acc) for each learner_id in other instances (Pfisterer et al., 2022; Binder et al., 2020). Values are reported as mean $\pm$ standard error, both computed over 10 independent runs with different random seeds. The learner_ids are sorted from best to worst max performance.
(a) Instance ID 1457
Learner ID Min Mean Max
ranger $0.014\pm 0.001$ $0.555\pm 0.004$ $0.832\pm 0.003$
svm $0.000\pm 0.000$ $0.211\pm 0.008$ $0.803\pm 0.003$
glmnet $0.007\pm 0.000$ $0.146\pm 0.002$ $0.657\pm 0.019$
rpart $003\pm 0.000 0.0$$063\pm 0.002$ $0.362\pm 0.006$
aknn $0.014\pm 0.000$ $0.132\pm 0.002$ $0.354\pm 0.003$
xgboost $008\pm 0.001 0.0$$046\pm 0.001$ $0.354\pm 0.017$
(b) Instance ID 1063
Learner ID Min Mean Max
svm $0.711\pm 0.006$ $0.816\pm 0.001$ $0.933\pm 0.005$
xgboost $0.524\pm 0.008$ $0.746\pm 0.001$ $0.902\pm 0.002$
rpart $0.733\pm 0.002$ $0.818\pm 0.001$ $0.897\pm 0.003$
ranger $0.727\pm 0.001$ $0.818\pm 0.000$ $0.874\pm 0.002$
glmnet $0.728\pm 0.003$ $0.802\pm 0.001$ $0.883\pm 0.002$
aknn $0.187\pm 0.002$ $0.649\pm 0.006$ $0.846\pm 0.001$
(c) Instance ID 1479
Learner ID Min Mean Max
aknn $0.449\pm 0.002$ $0.630\pm 0.003$ $0.987\pm 0.002$
svm $0.418\pm 0.003$ $0.554\pm 0.003$ $0.820\pm 0.009$
xgboost $0.406\pm 0.007$ $0.524\pm 0.001$ $0.652\pm 0.006$
glmnet $0.430\pm 0.002$ $0.513\pm 0.001$ $0.646\pm 0.006$
rpart $0.444\pm 0.002$ $0.498\pm 0.001$ $0.629\pm 0.008$
ranger $0.446\pm 0.005$ $0.513\pm 0.001$ $0.583\pm 0.005$
(d) Instance ID 15
Learner ID Min Mean Max
ranger $0.588\pm 0.003$ $0.905\pm 0.002$ $0.985\pm 0.000$
svm $0.776\pm 0.008$ $0.921\pm 0.001$ $0.980\pm 0.001$
aknn $0.312\pm 0.001$ $0.721\pm 0.006$ $0.972\pm 0.001$
rpart $0.648\pm 0.005$ $0.886\pm 0.002$ $0.963\pm 0.001$
xgboost $0.469\pm 0.007$ $0.714\pm 0.002$ $0.931\pm 0.003$
glmnet $0.646\pm 0.004$ $0.795\pm 0.001$ $0.929\pm 0.001$
(e) Instance ID 1468
Learner ID Min Mean Max
svm $0.049\pm 0.004$ $0.471\pm 0.010$ $0.968\pm 0.001$
ranger $0.087\pm 0.003$ $0.782\pm 0.004$ $0.943\pm 0.002$
aknn $0.252\pm 0.014$ $0.808\pm 0.002$ $0.934\pm 0.002$
glmnet $0.102\pm 0.002$ $0.404\pm 0.005$ $0.890\pm 0.003$
rpart $0.083\pm 0.001$ $0.309\pm 0.004$ $0.786\pm 0.007$
xgboost $0.080\pm 0.003$ $0.206\pm 0.001$ $0.744\pm 0.010$
(a) Instance ID 1457
(b) Instance ID 1063
(c) Instance ID 1479
(d) Instance ID 15
(e) Instance ID 1468
Figure 14. condPED-ANOVA HPIs on real-world problems from the YAHPO Gym rbv2_super scenario with other instances (Pfisterer et al., 2022; Binder et al., 2020). The bars denote the mean, and the error bars denote the standard error, both computed over ten independent runs with different random seeds.